# A Large-Scale Comprehensive Measurement of AI-Generated Code in Real-World Repositories

### 28 Mar 2026 v1

@@ Abstract.
Large language models (LLMs) are rapidly transforming software engineering by enabling developers to generate code ranging from small snippets to entire projects. As AI-generated code becomes increasingly integrated into real-world systems, understanding its characteristics and impact is critical. However, prior work primarily focuses on small-scale, controlled evaluations and lacks comprehensive analysis in real-world settings.
In this paper, we present a large-scale empirical study of AI-generated code in real-world repositories. We analyze both code-level metrics (e.g., complexity, structure, and defect-related indicators) and commit-level characteristics (e.g., commit size, frequency, and post-commit stability). To enable this study, we develop heuristic filter with LLM classification to identify AI-generated code and construct a large dataset.
Our results provide new insights into how AI-generated code differs from human-written code and how AI assistance influences development practices. These findings contribute to a deeper understanding of the practical implications of AI-assisted programming.

### 03 Apr 2026 v2

# A Large-Scale Empirical Study of AI-Generated Code in Real-World Repositories

@@ Abstract.
Large language models (LLMs) are increasingly used in software development, generating code that ranges from short snippets to substantial project components. As AI-generated code becomes more common in real-world repositories, it is important to understand how it differs from human-written code and how AI assistance may influence development practices. However, existing studies have largely relied on small-scale or controlled settings, leaving a limited understanding of AI-generated code in the wild.
In this work, we present a large-scale empirical study of AI-generated code collected from real-world repositories. We examine both code-level properties, including complexity, structural characteristics, and defect-related indicators, and commit-level characteristics, such as commit size, activity patterns, and post-commit evolution. To support this study, we develop a detection pipeline that combines heuristic filtering with LLM-based classification to identify AI-generated code and construct a large-scale dataset for analysis.
Our study provides a comprehensive view of the characteristics of AI-generated code in practice and highlights how AI-assisted development differs from conventional human-driven development. These findings contribute to a better understanding of the real-world impact of AI-assisted programming and offer an empirical basis for future research on AI-generated software.